OpenStudy (anonymous):

what equation has a graph parallel to the graph of 9x + 3y=22? This is not my strong point so i do not know what to do

OpenStudy (crabbyoldgamer):

Is this supposed to be multiple choice? 9x+3y=22 3y = -9x+22 y = -3x +22/3 Any line with a slope of -3 will be parallel to that line or the same line.

OpenStudy (campbell_st):

do you know the slope of your line as parallel lines have the same slope

OpenStudy (crabbyoldgamer):

In general, place a line equation in the form y = mx + b. It's called the slope intercept form. m is the slope, b is the y intercept. Then you can see which lines are parallel: they have the same slope (m).

OpenStudy (anonymous):

can you walk me through that ?

OpenStudy (crabbyoldgamer):

You want the equation in the form y = mx+b, where m is the slope. 9x + 3y = 22 You need the x component on the right side of =, so subtract 9x from both sides. 3y = -9x + 22 You want y alone, so divide both sides by 3. y = -3x + 22/3 y = mx + b: m = -3. Any line with a slope of -3 will either be the same line, or parallel to that line.

OpenStudy (anonymous):

Thanks!

OpenStudy (crabbyoldgamer):

Look up line slope intercept on purplemath.com for more explanation, or look in your book.

OpenStudy (anonymous):

kay bye